# artificial intelligence AoA angle-of-arrival

### v1

@@
Liqin Ding,  Gonzalo Seco-Granados, , Hyowon Kim,  Russ Whiton, Erik G. Ström,  Jonas Sjöberg, Henk Wymeersch This project has been supported in part by the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 101006664, and in part by the Catalan ICREA Academia Program. The authors would like to thank all partners within Hi-Drive for their cooperation and valuable contribution. This paper is extended from [1].L. Ding, E. G. Ström, J. Sjöberg, and H. Wymeersch are with the Department of Electrical Engineering, Chalmers University of Technology, Gothenburg, Sweden.G. Seco-Granados is with Department of Telecommunications and Systems Engineering, Universitat Autonoma de Barcelona, Barcelona, Spain.H. Kim is with the Department of Electronics Engineering, Chungnam National University, Daejeon 34134, South Korea.R. Whiton is with Volvo Car Corporation, Gothenburg, Sweden.
@@ Abstract
Ensuring positioning integrity amid faulty measurements is crucial for safety-critical applications, making receiver autonomous integrity monitoring (RAIM) indispensable. This paper introduces a Bayesian RAIM algorithm with a streamlined architecture for snapshot-type 3D cellular positioning. Unlike traditional frequentist-type RAIM algorithms, it computes the exact posterior probability density function (PDF) of the position vector as a Gaussian mixture (GM) model using efficient message passing along a factor graph. This Bayesian approach retains all crucial information from the measurements, eliminates the need to discard faulty measurements, and results in tighter protection levels (PLs) in 3D space and 1D/2D subspaces that meet target integrity risk (TIR) requirements. Numerical simulations demonstrate that the Bayesian RAIM algorithm significantly outperforms a baseline algorithm, achieving over $50\%$ PL reduction at a comparable computational cost.
@@ I Introduction
The advent of cellular networks brings promising enhancements in reliable positioning services, which are essential for applications like factory automation and autonomous driving [2, 3]. These sectors demand not only high accuracy and low latency but also stringent positioning integrity to ensure trustworthiness for downstream applications [4, 5]. While accuracy measures the closeness of the estimated position of a user equipment (UE) to its true position over a longer period and under nominal conditions, integrity assesses the real-time reliability of positioning estimates in the potential presence of measurement faults [6]. According to 3GPP, integrity guarantees for standalone radio positioning will be introduced in Release-18 and beyond [7, 8]. In this evolving landscape, receiver autonomous integrity monitoring (RAIM) [6, 9] is expected to play a crucial role.
Originally developed for global navigation satellite systems (GNSS) to meet the aviation industry’s stringent safety requirements [10, 11, 12, 13, 14, 15], RAIM has been adapted for various safety-critical positioning and navigation systems [16, 17, 18, 6, 9]. RAIM’s primary goal is to eliminate the effects of faulty measurements to achieve an accurate position estimate, and to compute an upper bound on the instantaneous position error, known as the protection level (PL). The PL ensures that the probability of the true position error exceeding this bound is less than the target integrity risk (TIR), thus quantifying integrity at a specified confidence level [6]. Traditional RAIM schemes achieve this through two steps: fault detection and exclusion (FDE) and error bounding. They rely on measurement redundancy and perform consistency checks either in the measurement domain (e.g., comparing measurement residuals [10, 11, 12, 13]) or in the position domain (e.g., using solution separation (SS) testing [14]) for FDE. These methods, which are frequentist in nature, tend to produce conservative PLs and thus lead to reduced availability [6].
The Bayesian RAIM framework, proposed relatively recently [19, 20], offers a robust alternative by assessing integrity based on the posterior probability distribution of the UE position, integrating all prior information and current measurements. Current research on Bayesian RAIM is limited, with most approaches relying on Monte Carlo methods such as particle filters [20, 21, 22] or Gibbs samplers [23] to approximate the posterior distribution. These methods, while powerful, can be computationally demanding and may not scale well with increased system complexity. In our prior work [1], we alternatively developed a Bayesian RAIM algorithm for time-of-arrival (ToA)-based cellular positioning in a simplified 1D setting using efficient message passing along a factor graph. The computed PLs are proven to be significantly tighter compared to traditional RAIM [14].
In this paper, we extend previous developments [1] to 3D positioning, addressing challenges in message passing and precise PL computation. Specifically, we focus on (i) accurately computing the weights of the GM models (ignored in [24, 25]), (ii) efficiently handling degenerate Gaussian densities in message passing; and (iii) precisely assessing the probability of a Gaussian distributed random vector within an arbitrary sphere in $\mathbb{R}^{n}$ for $n\geq 2$ for PL computation. Our main contributions are:
@@
  • Tight 2D/3D PL Computation: We develop a numerical integration-based method to accurately assess the probability of a Gaussian-distributed random vector residing within an arbitrary ellipsoid in $\mathbb{R}^{n}$. This advancement enables precise 2D/3D PL computations based on the probability density function (PDF) of the position, whether in Gaussian or Gaussian mixture (GM) model forms.
@@
The paper is organized as follows: Section II details the ToA-based 3D positioning problem and integrity monitoring, including the formal definition of the $n$-dimensional PL. Section III covers the Baseline RAIM algorithm. Section IV presents the message passing process for the Bayesian RAIM algorithm, and Section V describes the PL computation methods. The numerical study is in Section VI. Finally, conclusions are drawn in Section VII.
@@ II Problem Formulation
II-A System Model and Assumptions
We consider a single epoch of the positioning problem with $M$ time-synchronized base stations collaborating in downlink positioning of a UE. The UE’s four unknowns include its 3D position $\mathbf{x}_{\mathrm{u}}=(x_{\mathrm{u},1},x_{\mathrm{u},2},x_{\mathrm{u},3})^{\mathrm{T}}\in\mathbb{R}^{3}$ in the network coordinate reference system and its clock bias $x_{c}\in\mathbb{R}$. The positions of the $M$ BSs are known and expressed as $\mathbf{x}_{i}=(x_{i,1},x_{i,2},x_{i,3})^{\mathrm{T}}\in\mathbb{R}^{3}$ for $i=1,\ldots,M$. The BSs send coordinated positioning reference signals to the UE to estimate $M$ ToA for the line-of-sight (LoS) paths. These ToA convert to pseudoranges as ${d}_{i}=\|\mathbf{x}_{i}-\mathbf{x}_{\mathrm{u}}\|+x_{c}+b_{i}+n_{i}$, $i=1,\ldots,M$, where $n_{i}$ is measurement noise, and $b_{i}$ represents measurement bias caused by faults.
As usually assumed, an initial position estimate $\mathbf{x}_{\mathrm{u},0}$ is available. Linearizing $\|\mathbf{x}_{i}-\mathbf{x}_{\mathrm{u}}\|$ around $\mathbf{x}_{\mathrm{u},0}$ using a first-order Taylor expansion yields [26, 27, 28]
@@ II-A System Model and Assumptions
[…]
@@
We treat $\mathbf{x}\in\mathbb{R}^{4}$, $b_{i}$ and $n_{i}$ as realizations of the random vector/variable $\bm{X}=[X_{1},X_{2},X_{3},X_{4}]^{\mathrm{T}}$, $B_{i}$ and $N_{i}$, respectively. To indicate the measurement state of $y_{i}$, we introduce a latent random variable $\Lambda_{i}$ with realization $\lambda_{i}$, which follows the Bernoulli probability mass function (PMF) $p_{\Lambda_{i}}(\lambda_{i})=\theta_{i}^{\lambda_{i}}(1-\theta_{i})^{(1-\lambda_{i})}$ where $0<\theta_{i}\ll 1$. The value $\Lambda_{i}=0$ denotes a fault-free measurement with $B_{i}\sim\mathcal{N}(0,0)$, and $\Lambda_{i}=1$ indicates a faulty measurement with $B_{i}\sim\mathcal{N}(m_{b,i},\sigma_{b,i}^{2})$. Measurement noises are zero-mean Gaussian: $N_{i}\sim\mathcal{N}(0,\sigma_{n,i}^{2})$. The random variables $\{\Lambda_{i}\}$, $\{B_{i}\}$ and $\{N_{i}\}$ are considered independent within each set. The fault probability $\theta_{i}$, derived from previous system operations, is assumed to be known.
@@ II-B $n$-Dimensional Protection Level
The $n$-dimensional PL for the position estimate $\hat{\mathbf{x}}_{\mathrm{u}}$ in $\mathcal{L}(\tilde{\mathbf{v}}_{1})$ ($n=1$), $\mathcal{P}(\tilde{\mathbf{v}}_{1},\tilde{\mathbf{v}}_{2})$ ($n=2$), or $\mathbb{R}^{3}$ ($n=3$), for a TIR of $P_{\mathrm{TIR}}$, is the minimum distance $r$ such that the actual integrity risk (IR), $\mathrm{Pr}\left\{\|\mathbf{e}_{n\mathrm{D}}\|>r\right\}$, does not exceed $P_{\mathrm{TIR}}$
@@
For $n=2$ or $3$, given $\mathbf{e}_{n\mathrm{D}}=(e_{1},...,e_{n})^{\mathrm{T}}$ and weights [^1] Choosing $w_{i}=1/n$ for all $i$ is pragmatic when positioning error distributions are similar across coordinate directions. Adjusting these weights can refine the overestimate, leading to tighter results. Conversely, setting $w_{i}$ close to 00 can significantly increase $\mathrm{PL}_{1\mathrm{D},i}$ and consequently enlarge $\mathrm{PL}_{n\mathrm{D}}^{U}$. $\{w_{1},...,w_{n}\}$ such that $10<w_{i}<$ 1 for all $i$ and $\sum_{i=1}^{n}w_{i}=1$, if $\{\mathrm{PL}_{1\mathrm{D},1},...,\mathrm{PL}_{1\mathrm{D},n}\}$ are obtained such that
@@
Fig. 1 shows a scenario where the UE is equipped in a vehicle on a curved road in the horizontal-plane ($\mathrm{x}$-$\mathrm{y}$ plane). To ensure the vehicle remains within its lane, the position error perpendicular to the road ($\tilde{\mathbf{v}}_{\perp}$ direction) needs to stay within a margin at a confidence level of $1-P_{\mathrm{TIR}}$. As the vehicle navigates the curve, $\tilde{\mathbf{v}}_{\perp}$ changes. If $\tilde{\mathbf{v}}_{\perp}$ is known in real-time, the positioning system can compute $\mathrm{PL}_{1\mathrm{D}}(P_{\mathrm{TIR}})$ for $\mathcal{L}(\tilde{\mathbf{v}}_{\perp})$. Otherwise, it can provide an overestimated 2D PL $\mathrm{PL}_{2\mathrm{D}}^{U}=(\mathrm{PL}_{1\mathrm{D},1}^{2}+\mathrm{PL}_{1\mathrm{D},2}^{2})^{1/2}$, where $\mathrm{PL}_{1\mathrm{D},1}$ and $\mathrm{PL}_{1\mathrm{D},2}$ are 1D PLs computed for $\tilde{\mathbf{v}}_{1}=(1,0,0)^{\mathrm{T}}$ and $\tilde{\mathbf{v}}_{2}=(0,1,0)^{\mathrm{T}}$, with $w_{1}=w_{2}=0.5$. It is conceivable that by slightly decreasing $w_{1}$ and increasing $w_{2}$, the blue rectangle can become closer to a square, resulting in a tighter overestimate.
@@
PL requirements vary by application and must align with specific operational demands, see e.g. [9, Table II]. In Fig. 1, $\mathrm{PL}_{2D}^{U}$ might suffice for placing a vehicle on the correct road, suitable for automatic tolling, but it cannot ensure accurate lane positioning, which is crucial for automated driving.
III Baseline RAIM Algorithm
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Figure 1: Illustration of 2D PL computation on the horizontal plane.
In this process, the following components are involved: (i) all-in-view position estimation; (ii) fault mode identification; (iii) SS testing; and (iv) PL computations. The following sections provide detailed explanations of these components.
III-A All-in-View Position Estimation
III-B Fault Modes Identification
III-C Solution Separation Testing
$\hat{\mathbf{x}}^{(k)}=\mathsf{A}^{(k)}\mathbf{y}$ (9)
Remark 1 (Fault exclusion process).
When the SS test fails, the algorithm attempts to exclude the faulty measurements by reapplying the fault mode identification and SS testing to subsets of assumed fault-free measurements specified by the $N_{\mathrm{FM}}$ fault modes, ordered by decreasing probability of occurrence. This continues until the SS test passes or all fault modes are checked, indicating fault exclusion failure. Upon reaching fault mode $k$ in this process, the subset of measurements indexed by $\mathcal{I}_{k}$ are checked, and a new list of $N_{\mathrm{FM},k}=\sum_{j=1}^{|\mathcal{I}_{k}|-5}\binom{|\mathcal{I}_{k}|}{j}$ fault modes is identified. If the SS testing is successful, fault mode $k$ is considered true, and the algorithm returns the position estimate ($\hat{\mathbf{x}}_{\mathrm{u}}=\hat{\mathbf{x}}^{(k)}_{1:3}$) and PLs computed using these $|\mathcal{I}_{k}|$ measurements. If fault exclusion fails, the algorithm declares that the system cannot provide reliable positioning results (system unavailable).
III-D Protection Level Computation
A bisection search is used to find the smallest $r$ within a given error tolerance $r_{\mathrm{tol}}$. The detailed formulation process can be found in [14, Appendix H]. In the horizontal plane, $P_{\mathrm{TIR},\mathrm{H}}$ is evenly divided for the $\mathrm{x}$ and $\mathrm{y}$ directions (i.e., $w_{1}=w_{2}=0.5$ in (5)). For $n=1,2$, an overbound of the PL $\mathrm{PL}_{n}$ is given by the minimum $r_{n}$ that satisfies
Remark 2 (Complexity).
The Baseline RAIM algorithm has a variable computational complexity due to the varying number of SS tests required (from $1$ in the best case and $N_{\mathrm{FM}}$ in the worst case). The computational cost of a SS test mainly arises from (7)– (12), which involve matrix operations. The maximum size of the matrices is determined by the number of assumed fault-free measurements in the fault mode under examination, ranging between $M$ and $5$. To reduce complexity, fault modes with occurrence probabilities significantly lower than the TIR can be excluded from monitoring[14]. The PL computation using bisection search contributes little to the overall computational cost, which will be evident from numerical study. The bisection search is an iterative algorithm with variable number of iterations, denoted by $N_{\mathrm{it}}>0$. The complexity is $\mathcal{O}(N_{\mathrm{it}}N_{\mathrm{FM}})$ since both (13) and (14) involve approximately $N_{\mathrm{FM}}$ Q-function value computations.
IV Bayesian RAIM Part I: Message Passing
The linear Gaussian measurement model (2) allows a precise computation of the posterior probability distribution of $\bm{X}$, the random vector representing 3D position and clock bias, for each positioning epoch. This section details how the proposed Bayesian RAIM algorithm performs this computation through message passing along a factor graph.
IV-A Factor Graph Construction and Message Passing Schedule

### v2

@@
Liqin Ding, Gonzalo Seco-Granados, , Hyowon Kim, , Russ Whiton, Erik G. Ström, , Jonas Sjöberg, , Henk Wymeersch This project has been supported in part by the European Union’s Horizon 2020 research and innovation programme under grant agreement number 101006664, in part by the Swedish research council (VR) under grant number 2023-03821, in part by the Catalan Government under the ICREA Academia Program and grant 2021SGR00737, in part by the Spanish project PID2023-152820OB-I00 funded by MICIU/AEI/10.13039/501100011033 and ERDF/EU, in part by the National Research Foundation of Korea funded by the Ministry of Science and ICT under grant (RS-2025-25415939), and in part by the 2025 Daejeon RISE Project (DJR2025-12). The authors would like to thank all partners within Hi-Drive for their cooperation and valuable contribution. This paper is extended from [1].L. Ding was with the Department of Electrical Engineering, Chalmers University of Technology, Gothenburg, Sweden and is now with Ericsson Research, Ericsson AB, Lund, Sweden.G. Seco-Granados is with Department of Telecommunications and Systems Engineering, Universitat Autonoma de Barcelona, Bellaterra, Spain.H. Kim is with the Department of Electronics Engineering, Chungnam National University, Daejeon 34134, South Korea.R. Whiton was with Volvo Car Corporation, Gothenburg, Sweden and is now with European Space Agency, Noordwijk, Netherlands.E. G. Ström, J. Sjöberg, and H. Wymeersch are with the Department of Electrical Engineering, Chalmers University of Technology, Gothenburg, Sweden.
@@ Abstract
Ensuring positioning integrity amid faulty measurements is crucial for safety-critical applications, making receiver autonomous integrity monitoring (RAIM) indispensable. This paper introduces a Bayesian RAIM algorithm with a streamlined architecture for 3D cellular positioning. Unlike traditional frequentist-type RAIM algorithms, it computes the exact posterior probability density function (PDF) of the position vector as a Gaussian mixture (GM) model using efficient message passing along a factor graph. This Bayesian approach retains all crucial information from the measurements, eliminates the need to discard faulty measurements, and results in tighter protection levels (PLs) in 3D space and 1D/2D subspaces that meet target integrity risk (TIR) requirements. Numerical simulations demonstrate that the Bayesian RAIM algorithm significantly outperforms a baseline algorithm, achieving over $50\%$ PL reduction at a comparable computational cost.
@@ I Introduction
Ensuring integrity in positioning systems is vital for safety-critical applications such as autonomous driving, unmanned aerial vehicles, and industrial automation, where position errors can have catastrophic consequences [2, 3]. Integrity refers to the trustworthiness of the position information and the system’s ability to issue timely warnings when errors exceed acceptable thresholds [4, Chapter 7.5]. An important measure of integrity is integrity risk (IR), which is the probability that the error in the provided position information exceeds an acceptable tolerance without warning the user in a given period of time[2]. See [2, Section II] for additional integrity-related parameters and definitions.
With advanced cellular technologies like 5G and beyond, cellular network-based positioning has become a promising alternative or complement to global navigation satellite systems (GNSS), particularly in environments where GNSS is unreliable, such as indoors and dense urban areas [5, 6]. While cellular positioning has been extensively studied, with numerous algorithms developed to improve positioning accuracy [7, 8, 9], integrity issues remain underexplored. Robust methods are needed to ensure cellular positioning systems meet the stringent integrity requirements of safety-critical applications [10, 11, 12].
[…]
RAIM was originally developed for GNSS for aviation navigation and enables receivers to autonomously detect and mitigate faulty measurements [13, 14]. Traditional receiver autonomous integrity monitoring (RAIM) methods typically employ frequentist statistical approaches for fault detection and exclusion (FDE). These methods rely on measurement redundancy and perform consistency tests either in the measurement domain, such as analyzing measurement residuals [13, 14, 15, 16], or in the position domain using solution separation testing [17]. After FDE, the protection level (PL) is computed using error bounding methods to ensure that the integrity risk remains within the acceptable target integrity risk (TIR).
Applying traditional RAIM methods to cellular positioning systems presents significant challenges. On one hand, in typical cellular positioning scenarios, the number of available base stations is often limited, reducing the measurement redundancy needed for effective FDE [18]. Additionally, the challenging propagation environments, such as urban canyons and indoor settings, introduce multipath effects and non-line-of-sight (NLoS) conditions, complicating the signal measurements [19]. Inexpensive hardware may also introduce clock biases and other signal impairments [20]. As a result, faults are more prevalent in measurements using cellular signals, and these faults often differ in nature from those in GNSS. Traditional FDE methods may struggle to detect and exclude such faults effectively. On the other hand, traditional RAIM methods in GNSS compute separate PLs for horizontal and vertical directions, which is critical for aviation, while cellular positioning applications require PL computation in full 3D space. Land vehicles operate on varied terrains, and accurate vertical positioning is essential in environments with stacked highways, overpasses, or multi-level structures. UAVs and other airborne vehicles inherently require 3D positioning integrity [21]. Therefore, extending PL computation to 3D space, as well as to arbitrary 2D or 1D subspaces, is necessary for these applications. Moreover, excluding faulty measurements can discard valuable information, which is particularly detrimental when measurement redundancy is limited. This can lead to conservative PLs, reducing the availability of the positioning solution, as the system may declare the solution unsafe even when it is acceptable.
Figure 1: Illustration of the Bayesian RAIM framework. A long-term learning module uses hyper-priors to learn fault probabilities and bias models over time. For a single positioning epoch (as highlighted in the blue box and the focus of this paper), these fault probabilities and bias models are considered known. By incorporating dynamic models for both UE state and faults, the factor graph approach naturally extends to dynamic scenarios, enabling sequential belief updates for UE and fault tracking. This provides initial UE state estimates and updated fault priors for each positioning epoch.
These challenges can potentially be addressed by the Bayesian RAIM approach [22, 23], which integrates prior information and all measurements (including potentially faulty ones) into the positioning solution. By modeling faults probabilistically within the estimation process, this approach avoids fault exclusion and yields tighter PLs, thereby improving availability. However, existing Bayesian RAIM methods often depend on computationally intensive techniques such as particle filters or Monte Carlo sampling [23, 24, 25], making them unsuitable for real-time applications.
The Bayesian RAIM framework, after adapting our proposed assumptions and methodologies for time-of-arrival (ToA) based cellular positioning, is shown in Fig. 1. It supports long-term learning of measurement state and bias models through hyper-priors and allows integration of dynamic models for cross-epoch prediction. Focusing on a single epoch (the blue box in Figure 1), in previous work [1], we prototyped a Bayesian RAIM algorithm in a 1D setting using efficient message passing on factor graphs. The computed PLs were significantly tighter than those obtained by a traditional method [17]. In this paper, we extend this methodology to 3D positioning, addressing new challenges in message passing and PL computation in higher dimensions. Specifically, we focus on (i) accurate computation of Gaussian mixture (GM) model weights, which is often ignored in previous works [26, 27] but is crucial for precise posterior estimation; (ii) efficient handling of degenerate Gaussian densities in message passing, which is essential for computational efficiency and accuracy in multi-dimensional problems; (iii) precise computation of the probability that a Gaussian random vector lies within an arbitrary sphere in $\mathbb{R}^{n}$ for $n\geq 2$, which allows for application-specific integrity requirements under various scenarios. Our main technical contributions are as follows:
@@
  • Tight 2D/3D PL Computation: We develop a numerical integration-based method to accurately assess the probability of a Gaussian-distributed random vector residing within an arbitrary ellipsoid in $\mathbb{R}^{n}$. This advancement enables precise 2D/3D PL computations based on the probability density function (PDF) of the position, whether in Gaussian or GM model forms.
@@
The remainder of this paper is organized as follows. Section II presents the system assumptions and formally defines the $n$-dimensional PL. Section III describes the construction of the factor graph and details the message passing procedure used to derive the posterior position distribution. Section IV presents different methods for computing the PL. Section V presents a baseline RAIM algorithm that is employed for performance benchmarking. Numerical simulation results are presented in Section VI. Finally, Section VII offers the conclusions of this work.
@@ II Problem Formulation
II-A System Assumptions and the Bayesian RAIM Framework
We consider a cellular network where $M$ time-synchronized BSs collaborate in downlink positioning of a UE. The positions of the $M$ BSs in the network coordinate reference system: $\mathbf{x}_{i}=(x_{i,1},x_{i,2},x_{i,3})^{\mathrm{T}}\in\mathbb{R}^{3}$, $i=1,\ldots,M$, are known. In a single positioning epoch, the BSs send coordinated positioning reference signals to the UE to estimate $M$ ToA for the line-of-sight (LoS) paths. These ToA convert to pseudoranges as ${d}_{i}=\|\mathbf{x}_{i}-\mathbf{x}_{\mathrm{u}}\|+x_{c}+b_{i}+n_{i}$, $i=1,\ldots,M$, where $\mathbf{x}_{\mathrm{u}}=(x_{\mathrm{u},1},x_{\mathrm{u},2},x_{\mathrm{u},3})^{\mathrm{T}}\in\mathbb{R}^{3}$ and $x_{c}\in\mathbb{R}$ are the UE’s 3D position and clock bias, $n_{i}$ is measurement noise, and $b_{i}$ represents measurement bias caused by faults.
As is commonly assumed, an initial position estimate $\mathbf{x}_{\mathrm{u},0}$ is available. As illustrated in Fig. 1, it can be predicted based on the dynamic state model. Linearizing $\|\mathbf{x}_{i}-\mathbf{x}_{\mathrm{u}}\|$ around $\mathbf{x}_{\mathrm{u},0}$ using a first-order Taylor expansion yields [20, 28, 29]
@@ II-A System Assumptions and the Bayesian RAIM Framework
where $\mathbf{g}_{i}\triangleq(\mathbf{x}_{\mathrm{u},0}-\mathbf{x}_{i})/\|\mathbf{x}_{\mathrm{u},0}-\mathbf{x}_{i}\|$ is the unit vector from $\mathbf{x}_{\mathrm{u},0}$ to $\mathbf{x}_{i}$. Letting $\mathbf{h}_{i}=[\mathbf{g}_{i}^{\mathrm{T}}\;1]^{\mathrm{T}}$ and $\mathbf{x}=[\mathbf{x}_{\mathrm{u}}^{\mathrm{T}}\;x_{c}]^{\mathrm{T}}$, we establish a linear measurement model
@@
We treat $\mathbf{x}\in\mathbb{R}^{4}$, $b_{i}$ and $n_{i}$ as realizations of random vector/variable $\bm{X}=[X_{1},X_{2},X_{3},X_{4}]^{\mathrm{T}}$, $B_{i}$ and $N_{i}$, respectively. To indicate the measurement state of $y_{i}$, we introduce a latent random variable $\Lambda_{i}$ with realization $\lambda_{i}$, which follows the Bernoulli probability mass function (PMF) $p_{\Lambda_{i}}(\lambda_{i})=\theta_{i}^{\lambda_{i}}(1-\theta_{i})^{(1-\lambda_{i})}$ where $0<\theta_{i}\ll 1$. The value $\Lambda_{i}=0$ denotes a fault-free measurement with $B_{i}\sim\mathcal{N}(0,0)$, and $\Lambda_{i}=1$ indicates a faulty measurement with $B_{i}\sim\mathcal{N}(m_{b,i},\sigma_{b,i}^{2})$. Measurement noises are zero-mean Gaussian: $N_{i}\sim\mathcal{N}(0,\sigma_{n,i}^{2})$. The random variables $\{\Lambda_{i}\}$, $\{B_{i}\}$ and $\{N_{i}\}$ are considered independent within each set. By applying hyper-priors under the Bayesian framework, we can continually learn and refine both these models and their parameters over time. In the context of a single positioning epoch, we that assume these models and their parameters are known.
@@ II-B $n$-Dimensional Protection Level
The $n$-dimensional PL for the position estimate $\hat{\mathbf{x}}_{\mathrm{u}}$ in $\mathcal{L}(\tilde{\mathbf{v}}_{1})$ ($n=1$), $\mathcal{P}(\tilde{\mathbf{v}}_{1},\tilde{\mathbf{v}}_{2})$ ($n=2$), or $\mathbb{R}^{3}$ ($n=3$), for a TIR of $P_{\mathrm{TIR}}$, is the minimum distance $r$ such that the actual IR, $\mathrm{Pr}\left\{\|\mathbf{e}_{n\mathrm{D}}\|>r\right\}$, does not exceed $P_{\mathrm{TIR}}$
@@
For $n=2$ or $3$, given $\mathbf{e}_{n\mathrm{D}}=(e_{1},...,e_{n})^{\mathrm{T}}$ and weights [^1] Choosing $w_{i}=1/n$ for all $i$ is pragmatic when positioning error distributions are similar across coordinate directions. Adjusting these weights can refine the overestimate, leading to tighter results. Conversely, setting $w_{i}$ close to 0 can significantly increase $\mathrm{PL}_{1\mathrm{D},i}$ and consequently enlarge $\mathrm{PL}_{n\mathrm{D}}^{U}$. $\{w_{1},...,w_{n}\}$ such that $0<w_{i}<1$ for all $i$ and $\sum_{i=1}^{n}w_{i}=1$, if $\{\mathrm{PL}_{1\mathrm{D},1},...,\mathrm{PL}_{1\mathrm{D},n}\}$ are obtained such that
@@
[…]
@@
PL requirements vary by application and must align with specific operational demands, see e.g. [30, Table II]. In Fig. 2, $\mathrm{PL}_{2D}^{U}$ might suffice for placing a vehicle on the correct road, suitable for automatic tolling, but it cannot ensure accurate lane positioning, which is crucial for automated driving.
III Bayesian RAIM Part I: Message Passing
The linear Gaussian measurement model (2) allows a precise computation of the posterior probability distribution of $\bm{X}$, $p_{\bm{X}}^{\mathrm{pos}}(\mathbf{x})$, for each positioning epoch. This section describes the process for obtaining this distribution using message passing on a factor graph. The section concludes with several remarks, including a discussion of computational complexity and an explanation of why measurement exclusion is not recommended.
III-A Factor Graph Construction and Message Passing Schedule